Which of the following statements describes one part of completing the square for x² + 4x = 32?


a)Take the square root of 36 and add 2.
b)Take the square root of 32 and subtract 2.
c)Take the square root of 36 and subtract 2.

Answer by checkley79(3341)

SQRT36=6+2=8 ANS.
SQRT32=5.65-2=3.65 ANS.
SQRT36=6-2=34 ANS.